# The adelic closure of triangle groups

### v1

@@ 4.7. The $p$-adic image when $n=2p^{k}$
$0\rightarrow M\rightarrow\mathrm{SL}_{2}(\mathcal{O}/\mathfrak{p}^{d})\rightarrow\mathrm{SL}_{2}(\mathbf{F}_{p})\rightarrow 0$
By Lemma 4.5, the Frattini quotient of $M$ has rank $3$ over $\mathbf{F}_{p}$ and is identified with the image of $M$ in $\mathrm{SL}_{2}(\mathbf{F}_{p}[x]/x^{2})$. Let $N$ be the preimage of $\langle U\rangle$, so $N$ contains $M$ with index $p$. When $n=2p^{k}$, the image $G$ of $\Delta$ in $\mathrm{SL}_{2}(\mathcal{O}/\mathfrak{p}^{d})$ is a subgroup of $N$, since $U$ maps to $N$ by construction and $T$ maps to $M$. We have the following elements in $G\subset N$ and which are non-trivial in $N/M(\mathfrak{p}^{2})$:
(13) $V:=U^{(p+1)}\equiv$ $\ \left(\begin{matrix}1&0\\ 1&1\end{matrix}\right)\bmod\pi^{2},$
@@ 5. The $2$-adic image when $n=2^{k}$
    The image of $\Delta^{0}(2,\mathfrak{p})$ in $\Gamma^{0}(2,\mathfrak{p})/\Gamma^{0}(4,\mathfrak{p})$ is isomorphic to $\mathbf{Z}/2\mathbf{Z}\oplus(\mathcal{O}/2)^{2}$ given by $\widetilde{Q}$ in (15) together with ±$\pm I$.
@@
let us assume that $n\geq 8$. The image of $\Delta$ in $\mathrm{SL}_{2}(\mathcal{O}/2)$ is isomorphic to $D_{n/2}$. As follows from the computation of Section 3, the images of $T^{2}$, $U^{2}$, and $R^{d}$ with $d=n/2$ all lie in $\Gamma^{0}(2,\mathfrak{p})$, and all have the form (15). This proves that the image of $\Delta$ in $\Gamma^{0}(1,\mathfrak{p})/\Gamma^{0}(2,\mathfrak{p})$ is $D_{n/2}$, and (since matrices of the form (15) together with $-I$ are invariant under conjugation) that the image of $\Delta^{0}(2,\mathfrak{p})$ in $\Gamma^{0}(2,\mathfrak{p})/\Gamma^{0}(4,\mathfrak{p})$ is satisfies the required containment. Hence it suffices to show that the image contains all of this group. On the other hand, the images of $-I$,  $U^{2}$ and $T^{2}$ in this group in the basis of $\mathcal{O}/2\oplus\widetilde{Q}\simeq(\mathcal{O}/2)^{3}$ given by $(a,b,c)$ as in equation 14 have the form:
@@ 6. The invariants $\delta$ and $\kappa$
Let us now consider the invariants $\delta$ and $\kappa$ introduced in [McM24]. As previously mentioned, for any cusp $x\in\mathbf{P}^{1}(K)$, the invariant $\kappa(x)$ is defined to be 00 if $x$ is in the orbit of $\infty$ under $\Delta$ and $1$ otherwise. (In particular, if $n$ is odd and there is only one class of cusp, then $\kappa(x)$ always equals zero.) To prove that $\kappa$ is not a congruence invariant for some even $n$, it suffices to construct cusps $x$ adelically close to $\infty$ such that $\kappa(\infty)=1$. We phrase this more explicitly as follows. Suppose that
@@
Then $\gamma(0)=b/d$ is a cusp equivalent to 00, so to show that $\kappa$ is non-congruence it suffices to find, for any integer $N$, an element with $d\equiv 0\bmod N$, so $[b/d]=[\infty]$ in $\mathbf{P}^{1}(\mathcal{O}/N)$ but $\kappa(b/d)=1$.
The second invariant $\delta$ is defined as follows. Let $n^{\prime}=n$ if $n$ is odd and $n/2$ otherwise. (The integer $n^{\prime}$ was denoted by $m$ in [McM24] but that conflicts with our notation.) The image of $\Delta$ in $\mathrm{SL}_{2}(\mathcal{O}/2)$ is identified with $D_{n^{\prime}}=\langle r,t\rangle\subset\mathrm{Sym}(\mathbf{Z}/m)$, where $r(x)=x+1$ and $t(x)=-x$. The cusps of $\Delta$ are given by the coset $\Delta/\langle T\rangle$, and the map $\Delta\rightarrow\mathrm{SL}_{2}(\mathcal{O}/2)$ sending $T$ to $t\in D_{n^{\prime}}$ maps $\Delta/\langle T\rangle$ to $D_{n^{\prime}}/\langle t\rangle$, and then the latter is identified with $\mathbf{Z}/n^{\prime}\mathbf{Z}$ by sending $g$ to $g(0)$. In [McM24], one can further extend $\delta$ to the cusps equivalent to 00, but we have no need to discuss this here. Note that $\delta$ extends to a map on $\Delta$ itself by
$\delta(\gamma)=\delta(\gamma(\infty)).$
As noted in [McM24], the map $\delta\mod 2$ is a homomorphism but $\delta$ is not a homomorphism in general. The map $\delta$ on $\Delta$ clearly only depends on the congruence class of $\gamma\in\Delta$. However, as a map on cusps, this is no longer necessarily true. To show (for a particular even integer $n$) that $\delta$ is not a congruence invariant of the cusps, it suffices to find, for any integer $N$, a $\gamma\in\Delta$ such that:
@@
    The cusp $\gamma(\infty)=[a:c]$ is equal to $[1:0]\in\mathbf{P}^{1}(\mathcal{O}/N)$, equivalently, that $c\equiv 0\bmod N$.
The point is that the first condition implies that $\delta(\gamma(\infty))=i\bmod n/2$. Since (by [McM24, Cor 1.]) the invariant $2\delta$ is congruence, this should only be possible if $4|n$ and $i=n/4\bmod n/2$. Note that by the proof of [McM24, Theorem 1.13], it indeed suffices to find such a $\gamma$ for $N$ any power of $2$, although we don’t need to use this fact.
@@
The map $\delta\mod 2^{k-1}$ is a not a congruence invariant if $n=2^{k}m$ where $k\geq 2$ and $m>1$ is odd.
@@
In particular, since we can choose $b+c\in\mathcal{O}/2$ to be anything. we can ensure that $\gamma_{2}=\eta_{1}\gamma_{1}$ has a 00 in the first entry. Just as in the proof of Theorem 6.1, we proceed by induction to find the required element $\gamma_{\infty}$. ∎
@@
modulo any power of an odd prime $p$. On the other hand, by Lemma 6.2, there exists an element of the required form modulo any power of $2$, which proves the claim. But that means that the cusp $\gamma(\infty)=(a,c)\in\mathbf{P}^{1}(\mathcal{O}/N)$ maps to the image of 00 in $\mathbf{P}^{1}(\mathcal{O}/N)$, and so $\kappa$ is not a congruence invariant modulo $N$. Since this holds for any integer $N$, we are done. ∎

### v2

@@ 4.7. The $p$-adic image when $n=2p^{k}$
$0\rightarrow M\rightarrow\mathrm{SL}_{2}(\mathcal{O}/\mathfrak{p}^{d})\rightarrow\mathrm{SL}_{2}(\mathbf{F}_{p})\rightarrow 0.$
By Lemma 4.5, the Frattini quotient of $M$ has rank $3$ over $\mathbf{F}_{p}$ and is identified with the image of $M$ in $\mathrm{SL}_{2}(\mathbf{F}_{p}[x]/x^{2})$. Let $N$ be the preimage of $\langle U\rangle$, so $N$ contains $M$ with index $p$. When $n=2p^{k}$, the image $G$ of $\Delta$ in $\mathrm{SL}_{2}(\mathcal{O}/\mathfrak{p}^{d})$ is a subgroup of $N$, since $U$ maps to $N$ by construction and $T$ maps to $M$. We have the following elements in $G\subset N$ which are non-trivial in $N/M(\mathfrak{p}^{2})$:
(14) $V=U^{(p+1)}\equiv$ $\ \left(\begin{matrix}1&0\\ 1&1\end{matrix}\right)\bmod\pi^{2},$
@@ 5. The $2$-adic image when $n=2^{k}$
(16) $\left(\begin{matrix}b+c&b(\zeta+\zeta^{-1})\\ c&b+c\end{matrix}\right).$
@@
    The image of $\Delta^{0}(2,\mathfrak{p})$ in $\Gamma^{0}(2,\mathfrak{p})/\Gamma^{0}(4,\mathfrak{p})$ is isomorphic to $\mathbf{Z}/2\mathbf{Z}\oplus(\mathcal{O}/2)^{2}$ given by $\widetilde{Q}$ in (16) together with $\pm I$.
@@
Let us assume that $n\geq 8$. The image of $\Delta$ in $\mathrm{SL}_{2}(\mathcal{O}/2)$ is isomorphic to $D_{n/2}$. As follows from the computation of Section 3, the images of $T^{2}$, $U^{2}$, and $R^{d}$ with $d=n/2$ all lie in $\Gamma^{0}(2,\mathfrak{p})$, and all have the form (16). This proves that the image of $\Delta$ in $\Gamma^{0}(1,\mathfrak{p})/\Gamma^{0}(2,\mathfrak{p})$ is $D_{n/2}$, and (since matrices of the form (16) together with I-I are invariant under conjugation) that the image of $\Delta^{0}(2,\mathfrak{p})$ in $\Gamma^{0}(2,\mathfrak{p})/\Gamma^{0}(4,\mathfrak{p})$ is satisfies the required containment. Hence it suffices to show that the image contains all of this group. On the other hand, the images of I-I,  $U^{2}$ and $T^{2}$ in this group in the basis of $\mathcal{O}/2\oplus\widetilde{Q}\simeq(\mathcal{O}/2)^{3}$ given by $(a,b,c)$ as in equation 15 have the form:
@@ 6. The invariants $\delta$ and $\kappa$
Let us now consider the invariants $\delta$ and $\kappa$ introduced in [McM24]. As previously mentioned, for any cusp $x\in\mathbf{P}^{1}(K)$, the invariant $\kappa(x)$ is defined to be 0 if $x$ is in the orbit of $\infty$ under $\Delta$ and $1$ otherwise. (In particular, if $n$ is odd and there is only one class of cusp, then $\kappa(x)$ always equals zero.) To prove that $\kappa$ is not a congruence invariant for some even $n$, it suffices to construct cusps $x$ adelically close to $\infty$ such that $\kappa(\infty)=1$. We phrase this more explicitly as follows. Suppose that
@@
Then $\gamma(0)=b/d$ is a cusp equivalent to 0, so to show that $\kappa$ is non-congruence it suffices to find, for any integer $N$, an element with $d\equiv 0\bmod N$, so $b/d]=[\infty]$ in $\mathbf{P}^{1}(\mathcal{O}/N)$ but $\kappa(b/d)=1$.
The second invariant $\delta$ is defined as follows. Let $n^{\prime}=n$ if $n$ is odd and $n/2$ otherwise. (The integer $n^{\prime}$ was denoted by $m$ in [McM24] but that conflicts with our notation.) The image of $\Delta$ in $\mathrm{SL}_{2}(\mathcal{O}/2)$ is identified with $D_{n^{\prime}}=\langle r,t\rangle\subset\mathrm{Sym}(\mathbf{Z}/m)$, where $r(x)=x+1$ and t(x)=xt(x)=-x. The cusps of $\Delta$ equivalent to $\infty$ are given by the coset $\Delta/\langle T\rangle$, and the map $\Delta\rightarrow\mathrm{SL}_{2}(\mathcal{O}/2)$ sending $T$ to $t\in D_{n^{\prime}}$ maps $\Delta/\langle T\rangle$ to $D_{n^{\prime}}/\langle t\rangle$, and then the latter is identified with $\mathbf{Z}/n^{\prime}\mathbf{Z}$ by sending $g$ to $g(0)$; this is the definition of $\delta$ on the cusps equivalent to $\infty$. Note that $\delta$ can also be considered as a map on $\Delta$ itself by the formula
(19) $\delta(\gamma)=\delta(\gamma(\infty)).$
6.1. Some Reductions
In [McM24], the map $\delta$ can be extended to a function defined on all the cusps. The proof of Theorem 1.4 in this case can be reduced to the rest of Theorem 1.4 as follows. If $\delta$ is not congruence as a function on the cusps equivalent to $\infty$, it is certainly not congruence as an invariant on all of the cusps. Thus, since $\delta$ is a congruence invariant on all cusps when $n=2^{k}$ (by [McM24, Cor 1.5]), and since there is only one orbit of cusps when $n$ is odd, it suffices to consider the case when $2$ exactly divides $n$. If $n=4g+2$, it follows from Theorem [McM24, 3.2] that $\delta$ (as a function on all the cusps) is a congruence invariant if and only if $\kappa$ is congruence, and then the result follows from our claims concerning $\kappa$.
From this point onwards, we now regard $\delta$ as function on the cusps equivalent to $\infty$, or equivalently as a function on the group $\Delta$ as described in equation (19).
6.2. The strategy
As noted in [McM24], the map $\delta\mod 2$ is a homomorphism but $\delta$ is not a homomorphism in general. The map $\delta$ on $\Delta$ clearly only depends on the congruence class of $\gamma\in\Delta$. However, as a map on cusps, this is no longer necessarily true. To show (for a particular even integer $n$) that $\delta$ is not a congruence invariant of the cusps equivalent to $\infty$, it suffices to find, for any integer $N$, a $\gamma\in\Delta$ such that:
@@ 6.2. The strategy
    The cusp $\gamma(\infty)=[a:c]$ is equal to $1:0]\in\mathbf{P}^{1}(\mathcal{O}/N)$, equivalently, that $c\equiv 0\bmod N$.
The point is that the first condition implies that $\delta(\gamma(\infty))=i\bmod n/2$. Since (by [McM24, Cor 1.7]) the invariant $2\delta$ is congruence, this should only be possible if $4|n$ and $i=n/4\bmod n/2$. Note that by the proof of [McM24, Theorem 1.13], it indeed suffices to find such a $\gamma$ for $N$ any power of $2$, although we don’t need to use this fact.
@@
The map $\delta\mod 2^{k-1}$ is a not a congruence invariant on the cusps equivalent to $\infty$ if $n=2^{k}m$ where $k\geq 2$ and $m>1$ is odd.
@@
In particular, since we can choose $b+c\in\mathcal{O}/2$ to be anything. we can ensure that $\gamma_{2}=\eta_{1}\gamma_{1}$ has a 0 in the first entry. Just as in the proof of Theorem 6.3, we proceed by induction to find the required element $\gamma_{\infty}$. ∎
@@
modulo any power of an odd prime $p$. On the other hand, by Lemma 6.4, there exists an element of the required form modulo any power of $2$, which proves the claim. But that means that the cusp $\gamma(\infty)=(a,c)\in\mathbf{P}^{1}(\mathcal{O}/N)$ maps to the image of 0 in $\mathbf{P}^{1}(\mathcal{O}/N)$, and so $\kappa$ is not a congruence invariant modulo $N$. Since this holds for any integer $N$, we are done. ∎